OpenStudy (anonymous):

Graph the expression sec2 x - sec2 x sin2 x on your calculator. Determine what constant or single function is equivalent to the given expression.

OpenStudy (joannablackwelder):

-tan(x-45) in deg mode

OpenStudy (joannablackwelder):

I don't usually just give answers like that, but I just got it by trial and error.

OpenStudy (anonymous):

so tan^-1 (45degrees)

OpenStudy (joannablackwelder):

No, not inverse tan, negative tan

OpenStudy (anonymous):

oh okay and if I put that down it'll be the right answer?

OpenStudy (joannablackwelder):

It is the exact same graph on my calculator.

OpenStudy (anonymous):

oh okay awesome! thank you so so so so so so so so so so so much!!!!!

OpenStudy (joannablackwelder):

Glad to help :)

OpenStudy (anonymous):

I have one more more quick question on another problem

OpenStudy (joannablackwelder):

Ok

OpenStudy (anonymous):

sin (x + y) - sin (x - y) = 2 cos x sin y Is this saying that on the right side there are 2cos(x)'s and 2 sin(y)'s

OpenStudy (anonymous):

that's wrong

OpenStudy (anonymous):

$$\sec^2(x)-\sec^2(x)\sin^2(x)=\sec^2(x)(1-\sin^2(x))=\sec^2(x)\cos^2(x)=1$$given \(\sec(x)=\dfrac1{\cos x}\)

OpenStudy (anonymous):

so it's identical to the graph of \(y=1\)

OpenStudy (anonymous):

\(\sin(x+y)-\sin(x-y)=2\cos(x)\sin(y)\) the \(2\cos(x)\sin(y)\) means \(2\cdot\cos(x)\cdot\sin(y)\) i.e. multiply them all together

OpenStudy (joannablackwelder):

I did sec(2x)-sec(2x)sin(2x)

OpenStudy (anonymous):

why of y=1? wouldn't it be of sec(x) because that's what 1/cosx is the identity of?

OpenStudy (anonymous):

@help!!!!! \(\sec^2(x)\cos^2(x)=\dfrac1{\cos^2(x)}\cdot\cos^2(x)=1\)

OpenStudy (anonymous):

that would be correct, then, @JoannaBlackwelder, but the question explicitly states >Determine what constant or single function is equivalent to the given expression. suggesting it's actually \(\sec^2(x)-\sec^2(x)\cos^2(x)\). @help!!!!! should tell us which it is

OpenStudy (anonymous):

I can't because I have an online graphing calculator for this class that won't graph this expression

OpenStudy (joannablackwelder):

We want to know what the problem statement says.

OpenStudy (joannablackwelder):

Is it a double angle or a squared function?

OpenStudy (anonymous):

Graph the expression sec2 x - sec2 x sin2 x on your calculator. Determine what constant or single function is equivalent to the given expression. These are the directions I was given, word-for-word

OpenStudy (anonymous):

each of the twos are squares

OpenStudy (anonymous):

so which answer is correct?

OpenStudy (joannablackwelder):

Oh, ok. Then mine is wrong.

OpenStudy (anonymous):

so the other one is correct?

OpenStudy (joannablackwelder):

Right.

OpenStudy (anonymous):

how about the other problem Verify the identity. sin (x + y) - sin (x - y) = 2 cos x sin y

OpenStudy (joannablackwelder):

Looks like you will need to use sum and difference formulas.

OpenStudy (anonymous):

thats what I was thinking of but I'm sure quite how to use them

OpenStudy (anonymous):

@help!!!!! replace \(\sin(x+y)=\sin x\cos y+\cos x\sin y\) and \(\sin(x-y)=\sin x\cos y-\cos x\sin y\)

OpenStudy (anonymous):

$$\begin{align*}\sin(x+y)-\sin(x-y)&=(\sin x\cos y+\cos x\sin y)-(\sin x\cos y-\cos x\sin y)\\&=\sin x\cos y+\cos x\sin y-\sin x\cos y+\cos x\sin y\end{align*}$$

OpenStudy (anonymous):

try to simplify that... also my previous answer was correct...

OpenStudy (anonymous):

does it simplify to 2sinxcosy2sinycosx?

OpenStudy (anonymous):

and okay

OpenStudy (anonymous):

not quite... notice it's \(\sin x\cos y+\dots\color{red}-\sin x\cos y+\dots\)

OpenStudy (anonymous):

so the \(\sin x\cos y\) terms cancel out giving us just \(\cos x\sin y+\cos x\sin y\) -- can you combine like terms? hint: \(a + a = 2a\)

OpenStudy (anonymous):

ohhhhhhhhhhh okay I see now, so then after they cancel out I'm left with 2cosx siny

OpenStudy (anonymous):

exactly! :-)

OpenStudy (anonymous):

awesome you are the greatest thank you for clearing the confusion with the coefficients :-) goodnight!